OpenStudy (anonymous):

Determine the slope and the y-intercept of the graph of the equation. x + 10y = 1

OpenStudy (chaise):

x + 10y = 1 Let's put this into point-slope formula? x-1=-10y -0.1x+0.1=y y=-0.1x+0.1 Your y intercept is -0.1 and your slope is also 0.1

OpenStudy (anonymous):

y = (-1/10) x + 1/10 slope = -0.1 and y-intercept ix (0,0.1)

OpenStudy (anonymous):

are you familiar with y=mx+b form? if so rearrange the equation into this form (ie: [y=-(x/10)+1/10\] and now u can read of this equation ... giving: m (slope) = -1/10 and 'b' (the y-intercept) =1/10

OpenStudy (saifoo.khan):